OpenStudy (anonymous):

The volume of a gas V held at a constant temperature in a closed container varies inversely with its pressure P. If the volume of a gas is 1000 cubic centimeters when the pressure is 450 mm Hg, find the volume when the pressure is 250 mm Hg.

OpenStudy (anonymous):

555.56

OpenStudy (anonymous):

given that V varies inversely with pressure V= kP given V (1000) and P (45) find k (constant of proportionality) V=kP k=V/P =1000/450 k=2.222.... find V if the pressure is 250 V= 2.222(250) v=555.555...

OpenStudy (anonymous):

or approximately 555.56

OpenStudy (anonymous):

thats what I thought! thank you!!

OpenStudy (anonymous):

glad to help. .

OpenStudy (anonymous):

medalssss :-)

OpenStudy (anonymous):

somehow its 1800? not 555.56

OpenStudy (anonymous):

its 1000*450 / 250